Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News, Travel, Visitors (McMullin)

  • Overall status: Running this week a 'stable' version of M&C/data acquisition system; focus on 3-bit test during the day/RSRO at night.
  • Next week Monday is a holiday; Monday morning meetings will happen on Tuesday; our meeting will shift to the end of those meetings (i.e., sometime between 10 and 11 start).
  • New proposal cycle will be announced today; current (Oct 1) is the last on the trimester schedule; we'll discuss further as the details are revealed; some concern that it will be difficult to judge proposals with the uncertainty in the technical capabilities available so far in advance.
  • Key dates:
    • 13 Sep: Begin move to DnC configuration
    • 13-17 Sep: 74 MHz array testing campaign
    • 15 Sep: Call for Proposals
      • Note: Ku/X-wide will not be offered in main call (allowed for exploratory).
    • 01 Oct: Proposal deadline (A config; lowband for end of B config)
    • 04 Oct: Begin move to C configuration
    • 13-15 Oct: Phased Array testing campaign
  • Staff Travel:
    • Hayward: 10 Aug - 30 Aug
    • Chandler: 01 Sep - 15 Sep
    • DRAO: Carlson: 16 Jul - 19 Aug
    • Rupen: 11 Sep - 21 Sep
    • Sowinski: 20 Sep - 15 Oct
    • McMullin: 01-10 Oct (vacation/Chile)
    • van Moorsel: 11-15 Oct
    • Wrobel: 24 Oct-27 Oct
    • DRAO: Dougherty: 11 Oct - 12 Oct (SKA; ABQ)
  • Visitors:
    • Brent Carlson: 18 October - 25 October; second trip post SKA meeting TBD
    • DRAO: Dougherty: 12 Oct - 17 Oct (NRAO-DSOC)
    • Adam Deller/Joe Lazio: 13-15 Oct (phased array testing)

Hardware status and updates (Langley, Perley, Rupen)

  • Array (Perley/Langley):
    • Notes: EA02 IF:C spikes - confirmed moved to antenna EA03; EA11 IF:B; seeing phase jumps on some antennas (EA18, EA20, EA27) - under review; new L, Ku rxs coming; last Q band by middle of September.
      • Antennas with 3-bit samplers: EA15, EA22, EA27, EA28
    • Links:
  • WIDAR (Rupen):
    • Notes: Currently have 3 quadrants populated; not tested; Oct 1st for new switch to be installed (allowing all racks to be populated/enabled). Martin is away so no CBE work; added new CM version yesterday (used in 3-bit testing).

3-bit testing status

  • Notes:3-bit/8-bit comparisons indicate higher noise; signs of internally generated RFI; problems with individual samplers; RCP vs LCP different. More testing throughout the week.

Software status and updates (Butler)

  • Links:
  • Notes: Ran over the weekend with the Version 0.9 system with good effect; still seeing some missing scans (AO255) that need to be hunted down; some issues with TelCal solutions (AB1346). Over the weekend, the archive stopped writing data (disk partition filled; cleaner has been tuned to work more often); additional issues with the archive ingest process; will notify ecsv when fixed. Running this week with the revised CM (supporting both switched power and 3/8-bit correlations).
    • Discussion of averaging in CBE: BB status: CM has capability of specifying different integration times (not yet tested); CBE has the capability for averaging in both time/freq (also not tested); testing these facilities has been lower priority than other areas (e.g., basic OSRO/RSRO functionality with no missing scans). Planning for delivery date.
    • BB asked about the OBIT filler to get at data; several people using this. Not guaranteed to work but would be much faster ingest to AIPS given the lower number of transformations. Within NRAO, OBIT is installed natively.
  • CASA Update (new functionality available/issues/enhancements). In CASA active version to be reviewed: flagcmd task (Myers), initial image viewer scriptability (Ott).
  • AIPS Update (new functionality available/issues/enhancements). See attachments. See 0-6% polarization (some antennas - others are more than 100%); dramatic improvement in calibration when using spectral PCAL; some discussion of discontinuities at subband boundaries (possibly from prior gain calibration); allows per spw and per channel solutions (in AIPS can also do some spectral smoothing prior to solving - in between these).
    • Leonia noted that he had worked with Greg Taylor on FARS task confirming its operation; still in progress.

AIPS PCAL enhancements (Perley/Greisen)

Observing recommendations update (Ops)

  • Notes: Still assessing final updates (CM); no change until complete as current guidelines are 'safe'.
    • Have accepted some 30 min calibration SBs as backups/makeups (as discussed last week). This does not decrement their allocated times.
    • Note that CALIBRATE_BANDPASS does not trigger TelCal solutions; might want to include CALIBRATE_AMPLI, etc to get those if monitoring.
  • Link: https://staff.nrao.edu/wiki/bin/view/EVLA/RSROObservingPreparationGuidelines

Commissioning priorities for the week (McMullin, Chandler)

  • 1) RSRO/ECSO science
  • 2) 3-bit sampler testing
  • 3) OSRO/RSRO blockers (focused on missing data/data quality issues)
  • 4) CSV testing
  • 5) Demo science observing
  • 6) OSRO science observing

Commissioning Reports

Spectral Line: Ott: Doppler setting recommendations/status.

  • See: https://bugs.aoc.nrao.edu/browse/ECSV-32 for report. Discussion of conclusion versus sampling theory; further analysis will be discussed within the spectral line group before setting a public recommendation.

Spectral Line: Momjian: Zeeman calibration test.

Switched power update (Butler/Rupen)

  • Being produced. 'Concentrator is turned on most days; will run overnight tonight. Action: Point Steve at a data set (smallish). Meeting this week on the flow through of information from SDM to MS needed for calibration of switched power.

In preparation

Polarization: Multi-band leakage/PA (Myers, Marvil): Needed.

Commissioning Test Planning

  • Spectral Line: Dhawan: Frequency offset tests and gain stability (this week). Report next week.
  • Polarization: Antenna cross polarization (Perley; Need script.); LSCX on 3C147; Tuesday-Wednesday (0800-0900 LST); Delayed
  • Polarization: Taylor: Calibration test. Data taken - report next week.
  • Spectral Line: Bandpass stability,multi-band on 3C84 et al. (Goss/Chomiuk); TBS
  • Other:

RSRO/ECSO Reports

Project Band(s) Test Code Test Status Allocation Status (D) Notes JIRA
RSRO            
AA30 Q TRSR0003 - 0/51 Ready
1773293 (5.5 hr; 0630-0830)
RSRO-1
AB1345 L TRSR0004 21May10 23/40 In Progress;Needs SB RSRO-2
AB1346 K TRSR0005
TSPE0001
23Jun10 15/42 Ready
1859763 (3.5 hr; 1630-1730)
RSRO-3
AC982 CKKaQ TRSR0002 28Jul10 30/30 Complete; pending data review RSRO-4
AH1006 LCXKa TRSR0006 - 16/36 In Progress; Need SBs RSRO-5
AK726 Ka TRSR0007 - 9/29.5 In Progress; Need SBs RSRO-6
AL746 CKa TRSR0008 05Jul10 8.5/11 In Progress
1779817 (2 hr; 1000-1230)
RSRO-7
AM1014 C TRSR0009 22Jun10 0/0 C config RSRO-9
AS1015 C TRSR0010 21May10 4/4 Complete RSRO-11
AT374 C TRSR0011 25Jul10 0/0 C config RSRO-12
AV318 L TRSR0012 - 12/12 Complete; PI converted to OSRO RSRO-13
ECSO          
AW775 L TRSR0017 05Jul10 44/72 In Progress
1660006 (7)
-
AM1005 KKaQ TRSR0014 - 3/20 In Progress
1833848 (3 hr; 1830-1830)
-
AM1038 Q TRSR0019 21May10 5/10 Needs SB -
AM1040 Ka TRSR0019 21May10 10/10 Complete -
AC992 L TRSR0018 26Jun10 4/4 Complete -
AO255 Ka TRSR0016 - 24/40 In progress
1756969 (4)
-
AM1018 Ka TRSR0015 - 0/48 No SBs submitted -
AB1353 Ka TRSR0013 - 0/51 No SBs submitted -
  • Issues looking at the data due to archive problems. AB1346- identified problems with some antennas (10, 20, 27 in particular - 10 has 0 amplitude post/pre mid-term bandpass cal); action: quick writeup by Thursday for discussion with engineers et al. AM1038 first observation looks good (no missing or mislabeled scans); some issues with EA20 RCP (Ken noted there was a broken digitizer found on Monday). For AH1006, Volker noted that he had lost (wasn't able to calibrate) the first hour of data due to lost calibrator data. Will submit more SBs soon; same with AK726.
    • Current Residents
      • Laura Perez (POC: Claire Chandler); area(s): High frequency observing
      • Dave Green (POC: Sanjay Bhatnagar): area(s): algorithm tests for wideband imaging (currently away).
      • Volker Heesen (POC: Michael Rupen); area(s): polarimetry
      • Peter Hofner (POC: TBD); area(s): post-processing
      • Russ Taylor (POC: Rick Perley); area(s): polarimetry

Demo Science (Chandler/Condon): Project Codes: *TDEM*

    • Status: Begun; list requires prioritization; limited remaining time in D-config; Will try for IRC+10216 this week as possible.
    • Testing
Project Status TimeSorted ascending Comments SB
Crab nebula Complete 1x10 L-band OSRO1/OSRO2 TDEM0007_sb1524351 (10 hr: 0030-1030 LST)
M87 Ready (HF) 1x6 Ka TDEM0006_sb1470572 (not submitted; 6 hr: 0900-1000 LST)
M51 Ready (MR) 1x8 L band; needs corr test TDEM0008_sb1542082 (7.75 hr: 0830-1130 LST)
M82 Complete 1x2 L band TDEM0010_sb1660855 (2 hr: 0400-0930 LST)
Herc A Ready 1x5 CX bands TDEM0011_sb1729424 (5 hr; 1500-1700)
Cygnus A Ready 3x6 1x10 possible; 12 OSRO hours obtained Withdrawn
IRC+10216 Ready (HF) 33 K, Ka, Q spectral sweep; no recirculation; pending test TDEM0003_sb1654086 (30 min: 0730-1130 LST); RSRO 1 GHz, dual pol; 125 kHz
1525160 (3 hr;0730-1130); OSRO 8 MHz BW (HC7N/HC5N)
1542567 (3 hr; 0730-1130); OSRO 8 MHz BW (HC5N/C3N)
1542814 (3 hr; 0730-1130 LST); OSRO 8 MHz BW (C3N)
1543061 (3 hr; 0730-1130); OSRO 8 MHz BW (C4H)

IRC+10216 preliminary report (Claussen)

  • See: https://bugs.aoc.nrao.edu/browse/ECDATA-287; some strong RFI on short baselines in one baseband. Project has been modified to do a 1 MHz resolution K band sweep with OSRO follow-up. Begun the calibration; will have the other tracks by the weekend so hopefully will be able to show some lines next week.

Commissioning Group updates

User Support Update (van Moorsel)

AOB

Misc Actions

  • [McMullin]: Splatalogue use for RFI line list. In progress.
  • [McMullin/Frail]: %REDWrite up disposition of RR items.
  • [McMullin/Butler/Rupen/Owen]: Resolve discussion of requirements for DnC test of 4-band.
  • [McMullin]: Confirm restriction of RSRO scan length at 90 s
  • [Owen]: Outline issue with 90s scan length restriction.
  • [McMullin/Durand]: Clean up outdated MainSaver issues.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [Rupen]: X, Ku band subbands are inverted (IF 1 is labelled as IF4, IF2 as 3, IF3 as 2, IF 4 as 1). Noted that this had been tested but there were enough changes to warrant a re-do.
      • Med Priority [TBD]: Phase offsets within scans (phase jump at scan boundaries). R&L phase offset on some scans - probably L302 LO problem, needs tracking down. MR: Seen on EA07, EA25 (both RR and LL)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised.
      • Low Priority [TBD]: First scan often bad.
      • Low Priority [TBD]: Phase (and amp) wobbles; long history.

-- JosephMcMullin - 29 Aug 2010
Topic attachments
I Attachment Action Size Date Who Comment
HERCA.pol.epseps HERCA.pol.eps manage 5 MB 2010-09-02 - 17:28 UnknownUser Herc A calibration with/without spectral PCAL
PDfile.epseps PDfile.eps manage 307 K 2010-09-02 - 17:29 UnknownUser PCAL C-band results.
Topic revision: r8 - 2010-09-02, jmcmulli
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback